OpenStudy (anonymous):

A circle has a circumference of 25.12 cm. What is its area? Use 3.14 for π. A. 4 cm2 B. 12.56 cm2 C. 25.12 cm2 D. 50.24 cm2

OpenStudy (anonymous):

Is B the correct answer?

OpenStudy (anonymous):

I believe so

OpenStudy (jdoe0001):

the circumference of it is 25.12 the circumference formulat is \(2\pi r^2\)

OpenStudy (anonymous):

b is the answer

OpenStudy (jdoe0001):

so, that means 25.12 = \(2\pi r^2 \implies r^2=\cfrac{25.12}{2\pi}\)

OpenStudy (anonymous):

Thanks a lot!!!

OpenStudy (jdoe0001):

from the Area formula, you have Area = \(\pi r^2\) so, use THAT \(r^2\) in the Area formula, see what you get

OpenStudy (jdoe0001):

wait.... the circumference is just \(\pi \ \ no\ \ ^2\)

OpenStudy (jdoe0001):

so, \(2\pi r \implies r=\cfrac{25.12}{2\pi}\)